### v1

# One Step is Enough: Multi-Agent Reinforcement Learning based on One-Step Policy Optimization for Order Dispatch on Ride-Sharing Platforms

Abstract
On-demand ride-sharing platforms face the fundamental challenge of dynamically bundling passengers with diverse origins and destinations and matching them with vehicles in real time, all under significant uncertainty. Recently, Multi-Agent Reinforcement Learning (MARL) has emerged as a promising solution for this problem, leveraging decentralized learning to address the curse of dimensionality caused by the large number of agents in the ride-hailing market and the resulting expansive state and action spaces. However, conventional MARL-based ride-sharing approaches heavily rely on the accurate estimation of Q-values or V-values, which becomes problematic in large-scale, highly uncertain environments. Specifically, most of these approaches adopt an independent paradigm, exacerbating this issue, as each agent treats others as part of the environment, leading to unstable training and substantial estimation bias in value functions. To address these challenges, we propose two novel alternative methods that bypass value function estimation. First, we adapt Group Relative Policy Optimization (GRPO) to ride-sharing, replacing the Proximal Policy Optimization (PPO) baseline with the group average reward to eliminate critic estimation errors and reduce training bias. Second, inspired by GRPO’s full utilization of group reward information, we customize the PPO framework for ride-sharing platforms and show that, under a homogeneous fleet, the optimal policy can be trained using only one-step rewards—a method we term One-Step Policy Optimization (OSPO). Experiments on a real-world Manhattan ride-hailing dataset demonstrate that both GRPO and OSPO achieve superior performance across most scenarios, efficiently optimizing pickup times and the number of served orders using simple Multilayer Perceptron (MLP) networks. Furthermore, OSPO outperforms GRPO in all scenarios, attributed to its elimination of bias caused by the bounded simulation time of GRPO. Our code, trained models, and processed data are publicly available at the anonymous repository: https://github.com/RS2002/OSPO.
Introduction
The widespread adoption of on-demand ride-sharing platforms has fundamentally transformed urban transportation systems, offering scalable and sustainable mobility solutions for modern cities. By enabling multiple passengers with similar itineraries to share a single vehicle, ride-sharing platforms alleviate congestion, reduce vehicle kilometers traveled, and decrease urban fleet sizes, thereby addressing pressing challenges such as traffic, emissions, and inefficient resource utilization (Wang and Yang 2019; Santi et al. 2014). However, managing such platforms presents significant operational complexity. Passengers arrive randomly over time, each with unique origins and destinations, and the platform must dynamically determine not only how to bundle new orders together, but also how to assign each bundle to a suitable vehicle. These dispatching decisions must take into account the current routing of each vehicle, the destinations of onboard passengers, and the service quality requirements of all users. Critically, all decisions must be made in real time, under uncertainty regarding future demand and supply, making the order dispatch problem a central and challenging task for ride-sharing platforms.
Recently, reinforcement learning (RL) methods have shown great promise in addressing the order dispatch problem on ride-sharing platforms. Although ride-sharing is fundamentally a centralized task—requiring the platform to aggregate information from all drivers and orders to make globally optimal decisions—the massive state and action spaces encountered in realistic settings (which can reach millions or billions (Sivagnanam et al. 2024)) pose a serious curse of dimensionality, rendering single-agent RL approaches intractable. As a result, most recent works adopt the Multi-Agent Reinforcement Learning (MARL) framework, where each driver is treated as an individual agent, and a global controller coordinates the agents to optimize the overall system objectives. MARL methods are typically categorized into three types: Decentralized Training with Decentralized Execution (DTDE), Centralized Training with Decentralized Execution (CTDE), and Centralized Training with Centralized Execution (CTCE) (Jin et al. 2025). Most existing ride-sharing approaches follow the DTDE paradigm, since centralized methods still face scalability limitations similar to single-agent RL, and are thus difficult to apply in large-scale scenarios with hundreds or thousands of agents. However, DTDE methods rely on the independent assumption that each agent acts based only on its local information and treats other agents as part of the environment. This often leads to unstable training and poor cooperation among agents, as well as large estimation errors in Q-values and V-values (Hu, Feng, and Li 2025), ultimately resulting in suboptimal and inefficient learning.
To address this problem, we propose two multi-agent policy optimization methods that do not require estimating Q-values or V-values. First, we adapt Group Relative Policy Optimization (GRPO) (Shao et al. 2024) for the ride-sharing task, replacing the V-value baseline in Proximal Policy Optimization (PPO) (Schulman et al. 2017) with the group average reward. This modification significantly reduces the computational resources required for training and minimizes bias by eliminating neural network estimation errors associated with V-values and Q-values. Furthermore, inspired by the baseline replacement strategy in GRPO, we introduce a more efficient method called One-Step Policy Optimization (OSPO), which can be trained using only one-step rewards by leveraging the homogeneous characteristics of agents in the ride-sharing task. Both methods are validated through case studies using real-world ride-hailing data and demonstrate superior performance compared to state-of-the-art approaches. The main contributions of this paper are summarized as follows:
  • We introduce a novel ride-sharing framework based on GRPO. By replacing the baseline in PPO with group average reward, our method eliminates the need to estimate V-values. It successfully improves training performance and stability by removing the influence of estimation bias. To the best of our knowledge, this is the first successful application of GRPO in a non-LLM and MARL context, specifically in ride-sharing platforms.
  • Inspired by GRPO, we further propose a more efficient policy optimization method named OSPO, where we derive that, according to the agent homogeneity property in ride sharing task, the advantage function can be replaced by the group relative reward within only one step. To the best of our knowledge, it is the simplest and most efficient RL method for the ride-sharing task. Furthermore, we also specify the conditions under which OSPO can be applied in identical MARL scenarios.
  • We validate the proposed methods using a real-world ride-hailing dataset from Manhattan. Experimental results show that both GRPO and OSPO consistently outperform existing ride-sharing approaches as well as conventional Policy Gradient (PG) methods across most scenarios, primarily by optimizing pickup times and increasing the number of served orders. Notably, both methods maintain high efficiency even when implemented with a simple Multilayer Perceptron (MLP) network, with OSPO achieving the lowest GPU utilization among all methods. Furthermore, OSPO surpasses GRPO in performance, which can be attributed to its elimination of the bias introduced by the bounded simulation time present in GRPO.
Figure 1: The training process consists of four steps: (1) Calculate the matching probability for each driver-order pair; (2) Assign orders by maximizing the probability score; (3) Calculate the group advantage; (4) Train the network using PPO-style policy gradients.
Preliminary
Group Relative Policy Optimization (GRPO)
With the success and popularity of Reinforcement Learning from Human Feedback (RLHF), RL-based LLM post-training has garnered increasing attention. This has also led to the emergence of many new RL methods, such as Direct Preference Optimization (DPO) (Rafailov et al. 2023), Reinforce Leave-One-Out (RLOO) (Ahmadian et al. 2024), and GRPO (Shao et al. 2024). Among these, GRPO is noted for its simple format, high efficiency, and promising performance, which are core factors contributing to the success of DeepSeekMath (Shao et al. 2024) and DeepSeek-V3 (Liu et al. 2024).
Serving as a PG (Sutton et al. 1999) based method, the optimization objective of GRPO has a similar format to PPO(Schulman et al. 2017), expressed as:
$\text{J}_{GRPO}(\theta)=\mathbb{E}_{s,u\sim\pi_{\theta^{-}}}\left[\min\left\{\frac{\pi_{\theta}(s)}{\pi_{\theta^{-}}(s)}\hat{\text{A}}_{\pi_{\theta}}(s,u),\text{CLIP}\right.\right.$ (1)
$\left.\left(\frac{\pi_{\theta}(s)}{\pi_{\theta^{-}}(s)},1-\epsilon,1+\epsilon^{\prime}\right)\hat{\text{A}}_{\pi_{\theta}}(s,u)\right\}\Bigg{]}-\beta\text{KL}(\pi_{\theta}\|\pi_{ref})\ ,$
where $\pi$ denotes the policy, $\theta$ and $\theta^{-}$ represent the current network parameters and the parameters used for experience collection, respectively, $s$ and $u$ denote the state and action, $\pi_{ref}$ is the reference policy (i.e., the initial policy before post-training), serving as a normalization term to prevent the policy from diverging too much from the reference, $\hat{\text{A}}_{\pi_{\theta}}$ represents the advantage function of policy $\pi_{\theta}$, and $\beta,\epsilon,\epsilon^{\prime}$ are hyper-parameters We follow the improvement in (Yu et al. 2025), which suggests setting a higher value for $\epsilon^{\prime}$ than for $\epsilon$ to encourage exploration.
Aside from the KL divergence term, the core difference between GRPO and PPO lies in the definition of the advantage function. In PPO, the original advantage function is defined as:
$\text{A}_{\pi}(s_{t},u_{t})=\text{Q}_{\pi}(s_{t},u_{t})-\text{V}_{\pi}(s_{t})\ ,$ (2)
where the Q-value can be estimated using real trajectories, and the V-value can be estimated through a critic network. In contrast, GRPO proposes a simpler and more efficient advantage function given by:
$\hat{\text{A}}_{\pi}(s_{t},u_{t})=\sum_{\tau\in K_{t}}\gamma^{\tau}\frac{\text{R}(s_{t+\tau},u_{t+\tau})-\mu}{\sigma}\ ,$ (3)
[…]
Problem Setup for Ride-sharing Order Dispatch
In this paper, we consider the optimal order assignment task of an on-demand ride-sharing platform with $n$ homogeneous drivers (agents), each equipped with a vehicle of capacity $c$. At each time step $t$, there will be $m_{t}$ newly arrived orders from customers in different locations. The platform must dynamically determine not only how to bundle new orders together, but also how to assign each bundle to a suit- able vehicle. These dispatching decisions must take into account the current routing of each vehicle, the destinations of onboard passengers, and the service quality requirements of all users. Specifically, we model this process as a Multi-Agent Markov Decision Process (MAMDP) (Littman 1994), defined as $<n,S,U,P,R,\gamma>$, where $n$ represents the number of agents, and $\gamma$ represents the discount factor. The global state $S$ can be written as $S=(s^{o},s_{1},s_{2},\ldots,s_{n})$, where $s^{o}$ represents the global state (i.e. information of orders to be assigned), visible to all agents, and $s_{i}$ represents the state of agent $i$. In our ride-sharing task, we set $s^{o}$ to consist of the OD and arrival times of all unconfirmed orders, while $s_{i}$ consists of the current location, remaining capacity, the cumulative reward since the episode started, the average cumulative reward of all agents, and the information of all onboard orders, including their destination, estimated total delivery time, and remaining delivery time. The global action $U$ is the joint action of all agents, expressed as $U=(u_{1},u_{2},\ldots,u_{n})$. At time $t$, the action $u_{i}$ is a $w_{t}$-dimensional 0-1 vector, where $u_{i,j}=1$ indicates that order $j$ is assigned to agent $i$, and $w_{t}$ is the total number of orders to be assigned at time $t$. The transition probability function $P$ is not modeled explicitly in model-free RL methods. For the reward function $\text{r}(s_{i},u_{i})$, we leave the detailed design in appendix. Since ride-sharing is a fully cooperative task, the global reward can be expressed as the sum of rewards from each individual agent, written as $\text{R}(S,U)=\sum_{i=1}^{n}\text{r}(s_{i},u_{i})$.
Methodology
In this section, we first describe how we adapt GRPO for the multi-agent ride-sharing scenario, and analyze the theoretical foundations that support its applicability. We then introduce our proposed OSPO method and specify the conditions under which it can be effectively applied.
GRPO for Ride Sharing
The workflow of GRPO for ride sharing is shown in Fig. 1, following a design similar to previous SAC-based work (Enders et al. 2023). Firstly, we require the policy network $\pi_{\theta}$ to directly output the action probability of each agent, expressed as:
$p_{i,t}=\pi_{\theta}([s_{i,t},s^{o}_{t}])\ ,$ (4)
where $s_{i,t}$ and $s^{o}_{t}$ represent the state of agent $i$ and the global state at time $t$, respectively, and $p_{i,t}\in[0,1]^{w_{t}}$ represents the action probability (i.e., the probability of matching each order to agent $i$). Due to the homogeneity of the agents, we utilize a shared network among all agents, similar to approaches taken in previous works.
Considering the action constraint that an order cannot be assigned to multiple agents, we cannot simply choose the action independently for each agent. Instead, we view the action probabilities as scores and use bipartite matching to maximize the assignment score, following (Enders et al. 2023; Hu, Feng, and Li 2025):
$\max_{U_{t}}\sum_{i\in\mathcal{I}}u_{i,j,t}\cdot y_{i,j,t},$ (5a)
s.t. $\sum_{i\in\mathcal{I}}u_{i,j,t}\leq 1,\quad\forall j\in\mathcal{J}_{t},$ (5b)
$\sum_{j\in\mathcal{J}_{t}}u_{i,j,t}\leq 1,\quad\forall i\in\mathcal{I},$ (5c)
$u_{i,j,t}\in\{0,1\},\quad\forall i\in\mathcal{I},j\in\mathcal{J}_{t},$ (5d)
where $u_{i,j,t}$ indicates whether agent $i$ is assigned order $j$ at time $t$ (with 1 indicating assignment and 0 indicating no assignment), and $y_{i,j,t}$ denotes the score $p_{i,j,t}$ of agent $i$ choosing order $j$ at time $t$ (with $y_{i,j,t}=-\infty$ for all unavailable workers at time $t$). The set =$\mathcal{I}=\{1,2,\ldots,n\}$ represents the agents, and the set $\mathcal{J}_{t}=\{1,2,\ldots,w_{t}\}$ represents the orders to be assigned. Constraint (5b) ensures that an order can be assigned to at most one agent, while constraint (5c) guarantees that each agent is assigned at most one order. After that, we can obtain the reward according to Eq. 13 and then optimize the policy by maximizing Eq. 1.
To make GRPO work in our homogeneous cooperative MARL scenario, only two minor changes should be made to the original objective function in the SARL scenario. First, when calculating the advantage function, we use the rewards of all agents over all steps to compute the mean value $\mu$ and standard deviation $\sigma$ in Eq. 3. (For stability, we propose to ignore samples with a reward of 0 when calculating $\mu$ and $\sigma$. This is feasible since those no-assignment samples do not correspond to a real action and therefore cannot be used for training.) Second, for the KL normalization term in Eq. 1, since we do not have a reference policy like the LLM post-training, we propose to regularly track policy performance and retain the best checkpoint. This best policy serves as the reference policy, helping to ensure that the current policy does not deviate significantly from the best one. A detailed algorithm is provided in appendix.
Next, we analyze why our modification to GRPO is valid, allowing the transfer from SARL to MARL. For the advantage function (Eq. 3) in GRPO, we can express it as:
$\hat{\text{A}}_{\pi}(s_{t},u_{t})=\frac{1}{\sigma}\left(\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{t+\tau},u_{t+\tau})-\sum_{\tau\in K_{t}}\gamma^{\tau}\mu\right)\ .$ (6)
For the term $\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{t+\tau},u_{t+\tau})$, we can view it as an estimation of $\text{Q}_{\pi}(s_{t},u_{t})$, similar to what we do in PPO (Schulman et al. 2017). For the term $\sum_{\tau\in K_{t}}\gamma^{\tau}\mu$, we can write it as $\sum_{\tau\in K_{t}}\gamma^{\tau}(\mathbb{E}_{s,u\sim\pi}[\text{R}(s,u)]+\xi)$, where $\xi$ is the estimation error, which will reduce to 00 as the number of sampled trajectories increases (i.e., as the group becomes larger).
The original application scenario of GRPO is in LLM post-training, where the reward function is designed through a process supervision approach (Wang et al. 2023). An important feature of such a reward function is that each step reward has a similar range, leading to: $\mathbb{E}_{s_{t},u_{t}\sim\pi}[\text{R}(s_{t},u_{t})]=\mathbb{E}_{s,u\sim\pi}[\text{R}(s,u)]+\xi_{t}$, where $\xi_{t}$ is a small residual and $\mathbb{E}_{t}[\xi_{t}]=0$. Thus, we have:
$\sum_{\tau\in K_{t}}\gamma^{\tau}(\mathbb{E}_{s,u\sim\pi}[\text{R}(s,u)]+\xi)$ (7)
$=$ $\sum_{\tau\in K_{t}}\gamma^{\tau}(\mathbb{E}_{s_{\tau},u_{\tau}\sim\pi}[\text{R}(s_{\tau},u_{\tau})]-\xi_{\tau}+\xi)$
$=$ $\text{V}_{\pi}(s_{t})+\sum_{\tau\in K_{t}}\gamma^{\tau}(\xi-\xi_{\tau})\ .$
If we ignore the residual term, we obtain $\hat{\text{A}}_{\pi}(s_{t},u_{t})\approx\frac{1}{\sigma}(\text{Q}_{\pi}(s_{t},u_{t})-\text{V}_{\pi}(s_{t}))$. Compared to the advantage in PPO (Eq. 2), there is only one additional factor $\frac{1}{\sigma}$. According to (Kurin et al. 2022), this can be viewed as normalization for the reward function, contributing to reduced deviation and more stable training. Furthermore, based on the theory of policy gradients (Sutton et al. 1999), adding any baseline or positive factor does not influence the policy direction, regardless of the action. As a result, we consider the advantage of GRPO (Eq. 3) to be a valid replacement for the advantage in PPO (Eq. 2).
One Step Policy Optimization (OSPO)
We note that in the ride-sharability platform, all vehicles have identical capacity and are assigned to customers purely based on spatial proximity, without any additional attributes or preferences influencing the dispatch decisions. As a result, in the long run, the cumulative contribution or average earnings of each driver per unit time must converge to the same value. This is because the inherent randomness in order arrivals and assignments is experienced equally by all drivers over time. Therefore, it is reasonable to expect that different agents will exhibit similar value functions at each time step, regardless of their individual states. In this case, we have the following property:
$\text{V}_{\pi}(s_{1,t})+\epsilon_{1,t}=\text{V}_{\pi}(s_{2,t})+\epsilon_{2,t}=\cdots$ (8)
$=\text{V}_{\pi}(s_{n,t})+\epsilon_{n,t}=\mathbb{E}_{i\in\mathcal{I},s_{i,t}\sim\pi}[\text{V}_{\pi}(s_{i,t})]\ ,$
where $\epsilon_{i,t}$ is a small residual term, with $\mathbb{E}_{i\in\mathcal{I}}[\epsilon_{i,t}]=0$. As a result, when we ignore the residual terms, the different agents do not differ significantly from the group in GRPO, making it feasible in our scenario. Based on the property in Eq. 8, we can recalculate the advantage function as follows:
$\text{A}_{\pi}(s_{i,t},u_{i,t})=\text{Q}_{\pi}(s_{i,t},u_{i,t})-\text{V}_{\pi}(s_{i,t})$ (9)
$=$ $\text{r}(s_{i,t},u_{i,t})+\gamma\text{V}_{\pi}(s_{i,t+1})-\text{V}_{\pi}(s_{i,t})$
$=$ $\text{r}(s_{i,t},u_{i,t})+\gamma\mathbb{E}_{j\in\mathcal{I},s_{j,t+1}\sim\pi}[\text{V}_{\pi}(s_{j,t+1})]$
$-\mathbb{E}_{j\in\mathcal{I},s_{j,t}\sim\pi}[\text{V}_{\pi}(s_{j,t})]+\gamma\epsilon_{i,t+1}-\epsilon_{i,t}$
$=$ $\text{r}(s_{i,t},u_{i,t})+\gamma\mathbb{E}_{j\in\mathcal{J},s_{j,t+1}\sim\pi}[\text{V}_{\pi}(s_{j,t+1})]$
$-\mathbb{E}_{j\in\mathcal{I},s_{j,t},u_{j,t},s_{j,t+1}\sim\pi}[\text{r}(s_{j,t},u_{j,t})+\gamma\text{V}_{\pi}(s_{j+1,t})]$
$+\gamma\epsilon_{i,t+1}-\epsilon_{i,t}$
$=$ $\text{r}(s_{i,t},u_{i,t})-\mathbb{E}_{j\in\mathcal{I},s_{j,t},u_{j,t}\sim\pi}[\text{r}(s_{j,t},u_{j,t})]+\gamma\epsilon_{i,t+1}-\epsilon_{i,t}\ .$
If we perform a similar operation to GRPO, using the average reward to replace the expectation, applying reward normalization (Kurin et al. 2022), and ignoring the residual term, we arrive at the advantage function:
$\hat{\text{A}}_{\pi}(s_{i,t},u_{i,t})=\frac{\text{r}(s_{i,t},u_{i,t})-\mu}{\sigma}\ .$ (10)
This indicates that we can use only the one-step reward for policy optimization, which we term One Step Policy Optimization (OSPO). Even though the format may seem counter-intuitive, we provide an intuitive explanation. The validity of Eq. 10 is based on Eq. 8: the V-values remain similar among agents at each step due to the homogeneity in the ride-sharing system. For any arbitrary agent, regardless of its actions in the current step, it can achieve a similar V-value as others in the next step. Therefore, under the independent assumption, the goal of each agent is to maximize its single-step reward. In Eq. 8, removing the baseline and factor leaves only the single-step reward $\text{r}(s_{i,t},u_{i,t})$, which aligns with our intuition. Compared to GRPO, our OSPO further improves computational efficiency. Moreover, our method demonstrates greater robustness since it is not influenced by the discount factor.
Additionally, we propose two changes to the advantage function. First, we suggest directly using the mean value and standard deviation at each step $t$ to replace the values over the entire episode:
$\bar{\text{A}}_{\pi}(s_{i,t},u_{i,t})=\frac{\text{r}(s_{i,t},u_{i,t})-\mu_{t}}{\sigma_{t}}\ .$ (11)
This approach not only makes $\mu_{t}$ closer to $\mathbb{E}_{j\in\mathcal{I},s_{j,t},u_{j,t}\sim\pi}[\text{r}(s_{j,t},u_{j,t})]$ in Eq. 9, but also further improves computational efficiency, achieving true one-step computation. Second, during training, to better align with the property described in Eq. 8, we propose a penalty term $\Delta\delta_{t}$, where $\delta_{t}$ represents the standard deviation of the current cumulative reward for each agent, and $\Delta\delta_{t}=\delta_{t+1}-\delta_{t}$. This penalty is designed to reduce the utility gap (reflecting the V-value) between agents. Consequently, the final advantage function is expressed as:
$\tilde{\text{A}}(s_{i,t},u_{i,t})=\frac{\text{r}(s_{i,t},u_{i,t})-\mu_{t}}{\sigma_{t}}-\alpha\Delta\delta_{t}\ ,$ (12)
where $\alpha$ is a hyper-parameter. (Note that when Eq. 8 is satisfied, $\Delta\delta_{t}$ is close to 0. We also include the term $-\alpha\Delta\delta_{t}$ in the advantage function of GRPO.)
Experiment
Experiment Setup
To validate our method, we conduct a series of experiments using a real-world ride-hailing dataset from Manhattan, New York City (Taxi and Commission 2024). The data collected from 19:00 to 19:30 on July 17, 2024, is used as the training set, which includes 3,726 valid orders. The data from other periods of that day is used as the testing set.
For the simulation, we set the total number of drivers to 1,000, with each car having a capacity of 3 and a default speed of 60 km/h. The optimal routing is achieved using the OSRM simulator (Luxen and Vetter 2011). The experiments are conducted using the PyTorch framework (Paszke et al. 2019) on a workstation running Windows 11, equipped with an Intel(R) Core(TM) i7-14700KF processor and an NVIDIA RTX 4080 graphics card. Due to device limitations, we set each episode length to 30 minutes, with each step representing 1 minute, resulting in a running time of approximately 40 to 120 seconds for different methods per episode. We maximize the training episodes at 1,000 and evaluate model performance every 10 episodes.
To further illustrate the efficiency of our method, we choose a simple four-layer MLP as the policy network, with 128 units in each hidden layer and LeakyReLU (Maas et al. 2013) as the activation function. Considering the variable action space (since the number of orders changes each time), we set the network input as a single driver-order pair, ignoring the relationships between orders. Finally, a softmax layer is used to normalize the output probabilities.
During training, our models utilized approximately 3.8GB (OSPO) and 5.2GB (GRPO) of GPU memory, with a batch size of 256. For exploration, we add gradually decreasing random noise, employing the same Binary Symmetric Channel (BSC) noise as in (Hu, Feng, and Li 2025). For optimization, we use the Adam optimizer (Kingma 2014) with an initial learning rate of $10^{-4}$ and a decay rate of 0.99.
In the following, we first compare our GRPO and OSPO ride-sharing methods with previously popular methods of different types, including DTDE, CTDE, and CTCE. Next, we conduct a series of ablation studies to detail the effects of different modules in GRPO and OSPO, comparing them with similar policy gradient-based methods, including Independent PPO (IPPO) (De Witt et al. 2020), Multi-Agent PPO (MAPPO) (Yu et al. 2022), and Independent PG (IPG) (Sutton et al. 1999).
Comparative Experiment
Figure 2: Training Process: Each method is trained three times, and the curve is smoothed using Exponential Moving Average (EMA) with $\alpha=0.1$. The shaded area represents the range of fluctuations, while the solid line indicates the average value. (For delivery time and detour time, only completed orders are counted, as these metrics are uncertain for unfinished orders.)
Table 1: Model Performance on Testing Sets: Bold entries represent the best results.
Method Configuration Order Amount in Testing Sets Method Network GPU (GB) 2,850 3,114 3,577 3,910 DeepPool MLP (20K) 3.97 10,690.95±86.93 12,368.83±117.75 12,569.92±77.15 12,990.07±25.07 BMG-Q GAT (117K) 4.28 10,982.45±187.60 12,637.12±116.89 13,064.66±142.65 13,600.33±92.30 HIVES GRU (16M) 6.01 11,044.05±8.30 12,787.39±40.97 12,576.75±46.25 12,097.87±72.32 Enders et al. MLP (118K) 8.19 10,036.89±88.64 12,160.56±39.53 12,296.22±188.86 12,128.95±123.22 CEVD MLP (23K) 21.45 10,842.98±31.36 12,702.10±37.39 13,263.08±136.67 13,609.91±17.34 GRPO MLP (20K) 5.17 11,372.81±31.73 13,015.88±78.18 13,644.95±116.94 12,964.86±99.84 OSPO MLP (20K) 3.82 11,467.87±5.39 13,046.27±15.29 13,771.56±81.43 13,429.67±69.79
To illustrate the efficiency of our methods, we compare them with several representative ride-sharing approaches from various categories. A more detailed overview of previous work is provided in the appendix.
  • DeepPool (Al-Abbasi, Ghosh, and Aggarwal 2019) (DTDE): DeepPool is one of the earliest MARL-based ride-sharing works, based on IDDQN (Tan 1993; Van Hasselt, Guez, and Silver 2016) and using a CNN as the Q-network. To align it with our experimental scenario, we replace the CNN with an MLP similar to our methods (without the softmax).
  • BMG-Q (Hu, Feng, and Li 2025) (DTDE): To achieve better cooperation among agents, BMG-Q is also based on IDDQN but utilizes Graph Attention Network (GAT) (Velickovic et al. 2017) to capture the relationships among neighboring agents.
  • HIVES (Hao and Varakantham 2022) (CTDE): Based on QMIX (Rashid et al. 2020), HIVES tries addressing the curse of dimensionality problem by proposing a novel hierarchical mixing structure.
  • Enders et al. (Enders et al. 2023) (CTDE): Based on Multi-Agent Soft Actor Critc (MASAC) (Haarnoja et al. 2018), Enders et al. propose allowing agents to choose whether to accept assigned orders, which helps the system serve more valuable orders instead of treating all orders equally.
  • CEVD (Bose et al. 2023) (CTCE): CEVD modifies the Value Decomposition (VD) paradigm (Sunehag et al. 2018) from CTDE to CTCE. Specifically, it combines the Q-values of each agent with those of their neighbors to create a new type of Q-value, similar to the motivation behind BMG-Q.
For fairness, we modify the state space and reward function to align closely with those used in this paper. Additionally, we also choose a Greedy method as a baseline, which directly matches the order to the closest driver. The training process and testing results are shown in Fig. 2 and Table 1, respectively. In Fig. 2, we illustrate the cumulative reward alongside the number of orders served, as well as the average delivery time, detour time, pickup time, and confirmation time for each order. We observe that the GRPO and OSPO outperform the others primarily by achieving lower pickup and confirmation times, allowing them to serve more orders. In Table 1, it is noteworthy that the OSPO, which has the smallest network size and lowest GPU utilization, achieves the best performance across three testing scenarios. However, we also find that the GRPO and OSPO do not perform as well when the order amount increases to 3,910, which is the only scenario where the order amount exceeds 3,726 in the training set. Notably, the two methods that perform relatively well in this scenario are BMG-Q (Hu, Feng, and Li 2025) and CEVD (Bose et al. 2023). The primary distinction of these methods compared to others is their use of neighbor information when calculating Q-values. In Fig. 2, we also observe that they achieve lower detour times, suggesting better cooperation among agents through neighbor information. These results imply that incorporating more global or neighbor information may be more beneficial in on-pick scenarios. Moving forward, it would be advantageous to explore the model’s performance when the training set includes multiple order amount scenarios.
Ablation Study
In the ablation study, we primarily investigate the effects of:
  • Reward Normalization: We compare the effect of using single-step reward normalization ($\mu_{t},\sigma_{t}$ in Eq. 11) versus using the whole episode reward ($\mu,\sigma$ in Eq. 10) to study the trade-off between computational efficiency and performance.
  • Deviation Punishment: We examine the impact of including deviation punishment ($\Delta\delta_{t}$ in Eq. 12) to determine whether it helps to maintain property (ii) (Eq. 8).
  • KL Regulation Term: We assess the effect of using a KL regulation term ($\text{KL}(\pi_{\theta}\|\pi_{ref})$ in Eq. 1) to illustrate whether the historical best policy can prevent model divergence.
Figure 3: Result of Ablation Study
Since it seems unreasonable to use single-step reward normalization in GRPO, we omit this experiment. Additionally, we compare our methods with conventional policy optimization methods, including IPPO, MAPPO, and IPG. As shown in Fig. 3, we notice that the proposed GRPO and OSPO outperform all variant versions, illustrating the efficiency of the proposed modules. For the policy optimization methods, we observed that only IPPO demonstrates relatively good performance (even when compared to our proposed methods), while IPG struggles with high derivation and low sample efficiency. Additionally, MAPPO suffers from the curse of dimensionality in the centralized critic network, resulting in failure to converge.
Additionally, we observe that OSPO demonstrates overall better performance than GRPO. This is because, although the V-values among agents are generally similar, they have been maintained over a longer time period. In our ride-sharing scenario, the number of served orders may differ slightly among agents at each time step, but this variation can be considered negligible over the long term. Since our simulation lasts only 30 minutes, this difference becomes more pronounced when calculating the advantage (i.e., using the group average cumulative reward as a baseline introduces a slight bias). In contrast, OSPO exhibits greater robustness, as it relies solely on one-step rewards and does not encounter this issue.
Discussions
In this section, we analyze the current limitations and future directions of our work.
  • Application Constraints: As mentioned earlier, our method primarily relies on the property that the V-values of different agents remain similar at any time step (Eq. 8). This limits the application scenarios of OSPO to cooperative homogeneous environments. It cannot be used in cooperative scenarios where some agents must sacrifice their own interests to maximize the global reward. However, this condition can mostly be met in independent MARL, as each agent’s learning goal is to maximize its own long-term reward. In the next step, it is worthwhile to explore a wider range of possible application scenarios for OSPO.
  • Independent Assumption: One important reason for proposing OSPO is to reduce the estimation error of Q-values or V-values in independent MARL scenarios. However, the current training approach of our method still follows independent learning. The challenges of poor cooperation in independent MARL are yet to be addressed. Future work could explore combining OSPO with other methods, such as communication mechanisms or more powerful networks, similar to the approach taken by BMG-Q.
  • Fairness: Since OSPO is fundamentally based on the similar V-value property (Eq. 8), and we incorporate a punishment term (Eq. 12) to encourage the model to adhere to this property, it is important to explore whether this approach contributes to fairness. Recently, assignment fairness and discriminatory scenarios have garnered increasing attention in gig systems (Shi et al. 2021).
Conclusion
In this paper, we propose two novel MARL methods for ride-sharing that do not rely on estimating V-values or Q-values. First, we successfully adapt GRPO to the ride-sharing context, achieving promising performance by eliminating the estimation errors associated with the V-value network in PPO. Building on the insights from GRPO, we further introduce the OSPO method, which enables policy optimization using only single-step rewards. To the best of our knowledge, OSPO is the simplest and most efficient reinforcement learning method for order dispatch on ride-sharing platforms to date. Extensive experiments on real-world ride-sharing data demonstrate that our methods consistently outperform previous approaches in most scenarios, while requiring only a simple MLP network and exhibiting low GPU utilization.
Appendix
Reward Function Design
The reward function is shaped following previous work (Hu, Feng, and Li 2025):
$\text{r}(s_{i},u_{i})$ =$=\begin{cases}\beta_{1}+\beta_{2}p^{in}_{i}-\beta_{3}p^{out}_{i}-\beta_{4}\chi_{i}-\beta_{5}\rho_{i},|u_{i}|=1\\ 0,|u_{i}|=0\end{cases}$ (13)
where $\beta_{1}$ to $\beta_{5}$ are non-negative weights representing the platform’s valuation of each term, $p^{in}_{i}$ and $p^{out}_{i}$ represent the income from passengers and the payout to drivers, respectively. The variables $\chi_{i}$ and $\rho_{i}$ represent the number of en-route orders that will exceed their scheduled time and the additional travel time of all en-route orders when the assigned order is added to the scheduled route of agent $i$, respectively. This reward function is designed to comprehensively consider the interests of the platform, drivers, and passengers, mimicking the operation of a real-world ride-sharing platform. It is important to emphasize that $p^{in}_{i}$ and $p^{out}_{i}$ are calculated based on the order distance and the additional travel distance for the agent, respectively. When calculating travel time, we will utilize the Traveling Salesman Problem (TSP) to optimize the agent’s route.
Algorithm
The detailed algorithm of our GRPO/OSPO for ride sharing method is illustrated in Algorithm 1.
Algorithm 1 GRPO for Ride Sharing
0:  Policy network $\pi_{\theta}$ with shared parameters $\theta$
0:  Best policy checkpoint $\pi_{\phi}$ (initialized as $\pi_{\theta}$)
0:  Environment simulator, Reward function (Eq. 13)
1:  Initialize experience buffer $\mathcal{D}\leftarrow\emptyset$
2:  for episode $=1$ to $M$ do
3:     Reset environment, get initial states $\{s_{i,0}\}_{i=1}^{n}$, $s^{o}_{0}$
4:     for time step $t=0$ to $T$ do
5:        for each agent $i\in\mathcal{I}$ do
6:           Compute action probabilities (Eq. 4)
7:        end for
8:        Solve assignment problem via BiPartite matching (Eq. 5), getting action $U_{t}$
9:        Execute assignment $U_{t}$, observe rewards $\{r_{i,t+1}\}_{i=1}^{n}$
10:        Store transition ($(\{s_{i,t}\},s^{o}_{t},U_{t},\{r_{i,t+1}\},\{s_{i,t+1}\},s^{o}_{t+1})$ in $\mathcal{D}$
11:     end for
12:     Policy Optimization:
13:     for epoch $=1$ to $M$ do
14:        Sample batch $\mathcal{B}\sim\mathcal{D}$
15:        Compute group advantage $\hat{A}_{t}$
16:        Update policy parameters $\theta$ by maximizing Eq. 1
17:     end for
18:     if current policy $\pi_{\theta}$ outperforms $\pi_{\phi}$ then
19:        Update best checkpoint: $\phi\leftarrow\theta$
20:     end if
21:  end for
Related Work
In this section, we provide a brief overview of current ride-sharing methods. Starting from the Independent Double Deep Q-Learning (IDDQN)-based DeepPool (Al-Abbasi, Ghosh, and Aggarwal 2019), independent DTDE methods have become a mainstream paradigm, as the large number of agents, along with the extensive state and action spaces, severely hinders the development of centralized methods. Subsequent works have primarily aimed to enhance algorithms by fostering better cooperation among agents. For example, Graph Neural Networks (GNNs) have been widely explored to effectively utilize neighboring information (Hu, Feng, and Li 2025). Li et al. (Li et al. 2019) also propose using mean field approaches to capture more useful global information.
Additionally, some works attempt to transfer the successes of CTDE and CTCE to ride-sharing, where centralized critics can facilitate better cooperation among agents. However, these approaches face various shortcomings and challenges. Lima et al. (De Lima et al. 2020) first introduce QMIX, but only consider tasks involving a limited number of drivers in simulated grid scenarios, which presents a significant gap to real-world applications. Hao et al. (Hao and Varakantham 2022) explore hierarchical structures to address the curse of dimensionality in QMIX (Rashid et al. 2020), but the input dimensions of the mixture network still increase with the number of agents. Bose et al. (Bose et al. 2023) propose a Value Decomposition (VD) (Sunehag et al. 2018) based method; however, the credit assignment challenge becomes more complex in ride-sharing scenarios with hundreds of agents.
Lastly, there exists a series of works aimed at improving system efficiency by jointly considering relocation (Zhang et al. 2024), order bundling (Jiang, Xu, and Varakantham 2025), and multi-modal transportation (Hu, Dong, and Li 2025). Since these topics extend beyond the scope of this paper, we do not discuss them further here.
Nevertheless, most methods rely on precise value-function estimation (e.g., Q/V-networks), which introduces bias and instability due to approximation errors (Hu, Feng, and Li 2025). Our work addresses this gap by proposing GRPO/OSPO—eliminating value estimation entirely through group rewards—for stable and scalable coordination in homogeneous MARL systems.